a. Your timeline

b. Problem Type: FVA PMT

c. Solution:

$10,000 x FVA factor (6%/12, 5*12) = FVA answer

$10,000 x 69.77 = $143.33

A)Tim wants to go to travel around the world for the next 4 years. He believes he will need $36,000/year to live on. He wants to know how much money he must have in his account now to be able to fund his travels. Assume the money will earn 4%.

B) Refer to the problem. What would your answer be if the rate were 18%?
Answer #1

a) Amount required Every year = 36,000
Rate = 4%

PV of Amount = 36,000* PV Annuity Factor(4%,4) = 36000* (1-(1+4%)-4)/4% = 130676.23

b) Answer if rate = 18%
PV of Amount = 36000* PV Annuity Factor(18%,4) = 36000* (1-(1+18%)-4)/18% = 96842.22

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well